# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ A.7.2 HSIC independence testing
The exact same reasoning as the one presented for MMD holds for HSIC using $\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}$ instead of $\lambda_{1}\cdots\lambda_{d}$ (for details, see Schrab et al., 2022b, Theorem 3 which extends the results of Albert et al., 2022, Corollary 2, and Kim et al., 2022, Section 8.5), with the only difference being in the derivation of the quantile bound, which we present in details here. The aim is to prove that, with probability at least $21-\beta/$ 2, we have
${q}_{1-\alpha}\leavevmode\nobreak$ $\lesssim\leavevmode\nobreak\ \frac{1}{N\sqrt{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (119)
@@
$q_{1-\alpha}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \max\!\left(\frac{\Sigma}{N}\ln\!\left(\frac{1}{\alpha}\right),\frac{M}{N^{3/2}}\ln\!\left(\frac{1}{\alpha}\right)^{3/2}\right),$ (120)
@@
$M\leavevmode\nobreak\ \coloneqq\leavevmode\nobreak\ \max_{1\leq i,j,r,s\leq N}\big{|}k_{\lambda}(X_{i},X_{j})\ell_{\mu}(Y_{r},Y_{s})\big{|}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}$ (121)
@@
$\Sigma^{2}\leavevmode\nobreak\ \coloneqq\leavevmode\nobreak\ \!\left(\frac{1}{N^{2}}\sum_{1\leq i,j\leq N}k_{\lambda}(X_{i},X_{j})^{2}\right)\!\left(\frac{1}{N^{2}}\sum_{1\leq i,j\leq N}\ell_{\mu}(Y_{i},Y_{j})^{2}\right).$ (122)
@@
with probability at least $41-\beta/$ 4, and
$\frac{1}{N^{2}}\sum_{1\leq i,j\leq N}\ell_{\mu}(X_{i},X_{j})^{2}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{1}{N\mu\cdots\mu_{d_{\mathcal{Y}}}}+\frac{(N-1)}{N}\!\left(\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}\ell_{\mu}(X_{i},X_{j})^{2}\right)\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\mu\cdots\mu_{d_{\mathcal{Y}}}}\log\left(\frac{1}{\beta}\right)$ (124)
with probability at least $41-\beta/$ 4. We deduce that, with probability at least $21-\beta/$ 2, it holds
$\Sigma\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\sqrt{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}\log\left(\frac{1}{\beta}\right).$ (125)
with probability at least $21-\beta/$ 2. We conclude that
$q_{1-\alpha}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{\log(1/\alpha)\log(1/\beta)}{N\sqrt{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}\,\max\left\{1,\,\sqrt{\frac{\log(1/\alpha)}{N\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}\right\}$ (126)
@@
$\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2(d_{\mathcal{X}}+d_{\mathcal{Y}})/(4s+d_{\mathcal{X}}+d_{\mathcal{Y}})}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\log(1/\alpha)}{N},$ (127)
and so the quantile bound of Equation 126 becomes
$q_{1-\alpha}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{\log(1/\alpha)\log(1/\beta)}{N\sqrt{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}.$ (128)
@@
$\big{\|}\psi\big{\|}_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2s/(4s+d)}$ (129)
@@ A.7.3 KSD goodness-of-testing testing
${\bm{s}}_{P}(x)\leavevmode\nobreak\ =\leavevmode\nobreak\ -(\nu+d)\frac{x}{\nu+\|x\|_{2}^{2}}\qquad\textrm{ giving }\qquad\big{\|}{\bm{s}}_{P}(x)\big{\|}_{2}^{2}\leavevmode\nobreak\ =\leavevmode\nobreak\ (\nu+d)\frac{\|x\|^{2}_{2}}{\big{(}\nu+\|x\|_{2}^{2}\big{)}^{2}}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{\nu+d}{4\nu}$ (130)
@@
$\left|\frac{\partial}{\partial x_{i}}k_{\lambda}(x,y)\right|\leavevmode\nobreak$ $=\leavevmode\nobreak\ \left|\frac{2\beta(x_{i}-y_{i})}{\lambda^{2}+\|x-y\|_{2}^{2}}\right|k_{\lambda}(x,y)\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\lambda}\,k_{\lambda}(x,y),$ (132)
$\left|\frac{\partial}{\partial y_{i}}k_{\lambda}(x,y)\right|\leavevmode\nobreak$ $\lesssim\leavevmode\nobreak\ \frac{1}{\lambda}\,k_{\lambda}(x,y),$ (133)
$\left|\frac{\partial}{\partial x_{i}\partial y_{i}}k_{\lambda}(x,y)\right|\leavevmode\nobreak$ $\lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{2}}\,k_{\lambda}(x,y).$ (134)
@@
$\big{\|}\nabla_{\!1}k_{\lambda}(x,y)\big{\|}_{2}^{2}\leavevmode\nobreak$ $\lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{2d}}\,k_{\lambda}(x,y)^{2},$ (135)
$\big{\|}\nabla_{\!2}k_{\lambda}(x,y)\big{\|}_{2}^{2}\leavevmode\nobreak$ $\lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{2d}}\,k_{\lambda}(x,y)^{2},$ (136)
$\big{\|}\nabla_{\!1}^{\top}\!\big{(}\nabla_{\!2}k_{\lambda}(x,y)\big{)}\big{\|}_{2}^{2}\leavevmode\nobreak$ $\lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{4d}}k_{\lambda}(x,y)^{2},$ (137)
@@
$h_{\lambda}(x,y)^{2}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{4d}}k_{\lambda}(x,y)^{2}$ (139)
for all $x,y\in\mathbb{R}^{d}$. Equation 139 pinpoints the main difference between the MMD and KSD cases, the proof structure will be the same but the additional scaling in $\lambda$ will be affect the rate and needs to be kept track of.
@@
similar to the MMD expression of Equation 114 but with a different ${\bm{\psi}}$ function. Since the KSD is a score-based discrepancy, we will characterise departures from the null with the difference in scores multiplied by the data density
${\bm{\psi}}(x)\leavevmode\nobreak\ \coloneqq\leavevmode\nobreak\ \Big{(}\nabla\log p(x)-\nabla\log q(x)\Big{)}q(x).$ (141)
@@
The statistic concentration is the same as Equation 105 for the MMD case, that is
$\big{|}U_{0}-{\mathrm{Kdisc}}^{2}\big{|}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \sqrt{\frac{\sigma_{1}^{2}}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$ (144)
with probability at least $21-\beta/$ 2. However, the bound on $\sigma_{1}^{2}\coloneqq{\mathrm{var}_{Z}\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})\right]}\right)}$ differs. Ideally, we would like to upper bound $\sigma_{1}^{2}$ by $\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}=\int_{\mathbb{R}^{d}}\big{\|}{{\mathbb{E}}_{Y}\!\left[k(x,Y){\bm{\delta}}(Y)\right]}\big{\|}^{2}_{2}\,\mathrm{d}x$, similarly to Equation 106 for the MMD case. However, we can only show $\sigma_{1}^{2}\lesssim\int_{\mathbb{R}^{d}}{{\mathbb{E}}_{Y}\!\left[\|k(x,Y){\bm{\delta}}(Y)\|_{2}^{2}\right]}\,\mathrm{d}x$ which is not tight enough. Hence, we simply upper bound $\sigma_{1}^{2}$ in terms of $\lambda$, using Equation 139 we get
$\sigma_{1}^{2}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ {{\mathbb{E}}_{Z}\!\left[\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})\right]}\right)^{2}\right]}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ {{\mathbb{E}}_{Z,Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})^{2}\right]}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{4d}}{{\mathbb{E}}_{Z,Z^{\prime}}\!\left[k_{\lambda}(Z,Z^{\prime})^{2}\right]}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{5d}}.$ (145)
@@
$\big{|}U_{0}-{\mathrm{Kdisc}}^{2}\big{|}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{5d/2}}\,\sqrt{\frac{1}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$ (146)
with probability at least $21-\beta/$ 2. We note that without the bound in $\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}$, we cannot use the trick of Equation 112 to cancel the $\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}$ terms and get a rate in $N^{-1}$ instead of $N^{-1/2}$.
@@
Adapting the reasoning of Equations 107, 108, 109, 110 and 111 to keep track of the new bandwidth scaling of Equation 139, we get that Equation 108 becomes
$\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ \mathbb{E}\left[h_{\lambda}(Z,Z^{\prime})^{2}\right]+\sqrt{\frac{\widetilde{\sigma}_{1}^{2}}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{4d}}\,\mathbb{E}\left[k_{\lambda}(Z,Z^{\prime})^{2}\right]+\sqrt{\frac{1}{N\lambda^{10d}}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{5d}}+\frac{1}{\lambda^{5d}}\sqrt{\frac{1}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{5d}}\log\left(\frac{1}{\beta}\right)$ (147)
with probability at least $21-\beta/$ 2, with
$\widetilde{\sigma}_{1}^{2}\leavevmode\nobreak$ $\coloneqq\leavevmode\nobreak\ {\mathrm{var}_{Z}\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})^{2}\right]}\right)}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ {{\mathbb{E}}_{Z}\!\left[\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})^{2}\right]}\right)^{2}\right]}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{8d}}\,{{\mathbb{E}}_{Z}\!\left[\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[k_{\lambda}(Z,Z^{\prime})^{2}\right]}\right)^{2}\right]}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \frac{1}{\lambda^{10d}}.$ (148)
Finally, following the reasoning of Equations 107, 108, 109, 110 and 111, we obtain the quantile bound
${q}_{1-\alpha}\leavevmode\nobreak$ $\lesssim\leavevmode\nobreak\ \frac{1}{N\lambda^{5d/2}}\sqrt{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right)$ (149)
holding with probability at least $21-\beta/$ 2.
@@
Using Equations 146 and 149, the type II error can be controlled as
@@
${\mathrm{Kdisc}}^{2}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \frac{1}{\lambda^{5d/2}\sqrt{N}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (151)
@@
Using the expression of Equation 140, the power guaranteeing condition of Equation 151 becomes
$\big{\|}{\bm{\psi}}\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \big{\|}{\bm{\psi}}-S_{\lambda}{\bm{\psi}}\big{\|}_{L^{2}}^{2}-\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}+\frac{C}{\lambda^{5d/2}\sqrt{N}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (152)
The difference with the MMD case (e.g., Equation 115) is the rate $N^{-1/2}$ instead of $N^{-1}$, the extra term $-\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}$, and the rate in $\lambda$. Here, we simply bound $-\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}$ by 0 (instead of being able to use this term to improve the rate in $N$ as in the MMD case), we get
$\big{\|}{\bm{\psi}}\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \big{\|}{\bm{\psi}}-S_{\lambda}{\bm{\psi}}\big{\|}_{L^{2}}^{2}+\frac{C}{\lambda^{5d/2}\sqrt{N}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (153)
@@
$\int_{\mathbb{R}^{d}}\big{\|}\xi\big{\|}^{2s}_{2}\,\big{\|}\widehat{\bm{\psi}}(\xi)\big{\|}^{2}_{2}\,\mathrm{d}\xi\leavevmode\nobreak\ \leq\leavevmode\nobreak\ (2\pi)^{d}$ (154)

### v2

@@ A.7.2 HSIC independence testing
The exact same reasoning as the one presented for MMD holds for HSIC using $\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}$ instead of $\lambda_{1}\cdots\lambda_{d}$ (for details, see Schrab et al., 2022b, Theorem 3 which extends the results of Albert et al., 2022, Corollary 2, and Kim et al., 2022, Section 8.5), with the only difference being in the derivation of the quantile bound, which we present in details here. The aim is to prove that, with probability at least $1-\beta/2$, we have
${q}_{1-\alpha}\penalty 10000\$$\lesssim\penalty 10000\ \frac{1}{N\sqrt{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (119)
@@
$q_{1-\alpha}\penalty 10000\ \lesssim\penalty 10000\ \max\!\left(\frac{\Sigma}{N}\ln\!\left(\frac{1}{\alpha}\right),\frac{M}{N^{3/2}}\ln\!\left(\frac{1}{\alpha}\right)^{3/2}\right),$ (120)
@@
$M\penalty 10000\ \coloneqq\penalty 10000\ \max_{1\leq i,j,r,s\leq N}\big|k_{\lambda}(X_{i},X_{j})\ell_{\mu}(Y_{r},Y_{s})\big|\penalty 10000\ \lesssim\penalty 10000\ \frac{1}{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}$ (121)
@@
$\Sigma^{2}\penalty 10000\ \coloneqq\penalty 10000\ \!\left(\frac{1}{N^{2}}\sum_{1\leq i,j\leq N}k_{\lambda}(X_{i},X_{j})^{2}\right)\!\left(\frac{1}{N^{2}}\sum_{1\leq i,j\leq N}\ell_{\mu}(Y_{i},Y_{j})^{2}\right).$ (122)
@@
with probability at least $1-\beta/4$, and
$\frac{1}{N^{2}}\sum_{1\leq i,j\leq N}\ell_{\mu}(X_{i},X_{j})^{2}\penalty 10000\$$=\penalty 10000\ \frac{1}{N\mu\cdots\mu_{d_{\mathcal{Y}}}}+\frac{(N-1)}{N}\!\left(\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}\ell_{\mu}(X_{i},X_{j})^{2}\right)\penalty 10000\ \lesssim\penalty 10000\ \frac{1}{\mu\cdots\mu_{d_{\mathcal{Y}}}}\log\left(\frac{1}{\beta}\right)$ (124)
with probability at least $1-\beta/4$. We deduce that, with probability at least $1-\beta/2$, it holds
$\Sigma\penalty 10000\ \lesssim\penalty 10000\ \frac{1}{\sqrt{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}\log\left(\frac{1}{\beta}\right).$ (125)
with probability at least $1-\beta/2$. We conclude that
$q_{1-\alpha}\penalty 10000\ \lesssim\penalty 10000\ \frac{\log(1/\alpha)\log(1/\beta)}{N\sqrt{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}\,\max\left\{1,\,\sqrt{\frac{\log(1/\alpha)}{N\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}\right\}$ (126)
@@
$\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}\penalty 10000\ =\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2(d_{\mathcal{X}}+d_{\mathcal{Y}})/(4s+d_{\mathcal{X}}+d_{\mathcal{Y}})}\penalty 10000\ \geq\penalty 10000\ \frac{\log(1/\alpha)}{N},$ (127)
and so the quantile bound of Equation˜126 becomes
$q_{1-\alpha}\penalty 10000\ \lesssim\penalty 10000\ \frac{\log(1/\alpha)\log(1/\beta)}{N\sqrt{\lambda_{1}\cdots\lambda_{d_{\mathcal{X}}}\mu\cdots\mu_{d_{\mathcal{Y}}}}}.$ (128)
@@
$\big\|\psi\big\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2s/(4s+d)}$ (129)
@@ A.7.3 KSD goodness-of-testing testing
${\bm{s}}_{P}(x)\penalty 10000\ =\penalty 10000\ -(\nu+d)\frac{x}{\nu+\|x\|_{2}^{2}}\qquad\textrm{ giving }\qquad\big\|{\bm{s}}_{P}(x)\big\|_{2}^{2}\penalty 10000\ =\penalty 10000\ (\nu+d)\frac{\|x\|^{2}_{2}}{\big(\nu+\|x\|_{2}^{2}\big)^{2}}\penalty 10000\ \leq\penalty 10000\ \frac{\nu+d}{4\nu}$ (130)
@@
$\left|\frac{\partial}{\partial x_{i}}k_{\lambda}(x,y)\right|\penalty 10000\$$=\penalty 10000\ \left|\frac{2\beta(x_{i}-y_{i})}{\lambda^{2}+\|x-y\|_{2}^{2}}\right|k_{\lambda}(x,y)\penalty 10000\ \lesssim\penalty 10000\ \frac{1}{\lambda}\,k_{\lambda}(x,y),$ (132)
$\left|\frac{\partial}{\partial y_{i}}k_{\lambda}(x,y)\right|\penalty 10000\$$\lesssim\penalty 10000\ \frac{1}{\lambda}\,k_{\lambda}(x,y),$ (133)
$\left|\frac{\partial}{\partial x_{i}\partial y_{i}}k_{\lambda}(x,y)\right|\penalty 10000\$$\lesssim\penalty 10000\ \frac{1}{\lambda^{2}}\,k_{\lambda}(x,y).$ (134)
@@
$\big\|\nabla_{\!1}k_{\lambda}(x,y)\big\|_{2}^{2}\penalty 10000\$$\lesssim\penalty 10000\ \frac{1}{\lambda^{2d}}\,k_{\lambda}(x,y)^{2},$ (135)
$\big\|\nabla_{\!2}k_{\lambda}(x,y)\big\|_{2}^{2}\penalty 10000\$$\lesssim\penalty 10000\ \frac{1}{\lambda^{2d}}\,k_{\lambda}(x,y)^{2},$ (136)
$\big\|\nabla_{\!1}^{\top}\!\big(\nabla_{\!2}k_{\lambda}(x,y)\big)\big\|_{2}^{2}\penalty 10000\$$\lesssim\penalty 10000\ \frac{1}{\lambda^{4d}}k_{\lambda}(x,y)^{2},$ (137)
@@
$h_{\lambda}(x,y)^{2}\penalty 10000\ \lesssim\penalty 10000\ \frac{1}{\lambda^{4d}}k_{\lambda}(x,y)^{2}$ (139)
for all $x,y\in\mathbb{R}^{d}$. Equation˜139 pinpoints the main difference between the MMD and KSD cases, the proof structure will be the same but the additional scaling in $\lambda$ will be affect the rate and needs to be kept track of.
@@
similar to the MMD expression of Equation˜114 but with a different $\bm{\psi}$} function. Since the KSD is a score-based discrepancy, we will characterise departures from the null with the difference in scores multiplied by the data density
${\bm{\psi}}(x)\penalty 10000\ \coloneqq\penalty 10000\ \Big(\nabla\log p(x)-\nabla\log q(x)\Big)q(x).$ (141)
@@
The statistic concentration is the same as Equation˜105 for the MMD case, that is
$\big|U_{0}-{\mathrm{Kdisc}}^{2}\big|\penalty 10000\ \lesssim\penalty 10000\ \sqrt{\frac{\sigma_{1}^{2}}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$ (144)
with probability at least $1-\beta/2$. However, the bound on $\sigma_{1}^{2}\coloneqq{\mathrm{var}_{Z}\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})\right]}\right)}$ differs. Ideally, we would like to upper bound $\sigma_{1}^{2}$ by $\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}=\int_{\mathbb{R}^{d}}\big\|{{\mathbb{E}}_{Y}\!\left[k(x,Y){\bm{\delta}}(Y)\right]}\big\|^{2}_{2}\,\mathrm{d}x$, similarly to Equation˜106 for the MMD case. However, we can only show $\sigma_{1}^{2}\lesssim\int_{\mathbb{R}^{d}}{{\mathbb{E}}_{Y}\!\left[\|k(x,Y){\bm{\delta}}(Y)\|_{2}^{2}\right]}\,\mathrm{d}x$ which is not tight enough. Hence, we simply upper bound $\sigma_{1}^{2}$ in terms of $\lambda$, using Equation˜139 we get
$\sigma_{1}^{2}\penalty 10000\ \lesssim\penalty 10000\ {{\mathbb{E}}_{Z}\!\left[\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})\right]}\right)^{2}\right]}\penalty 10000\ \lesssim\penalty 10000\ {{\mathbb{E}}_{Z,Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})^{2}\right]}\penalty 10000\ \lesssim\penalty 10000\ \frac{1}{\lambda^{4d}}{{\mathbb{E}}_{Z,Z^{\prime}}\!\left[k_{\lambda}(Z,Z^{\prime})^{2}\right]}\penalty 10000\ \lesssim\penalty 10000\ \frac{1}{\lambda^{5d}}.$ (145)
@@
$\big|U_{0}-{\mathrm{Kdisc}}^{2}\big|\penalty 10000\ \lesssim\penalty 10000\ \frac{1}{\lambda^{5d/2}}\,\sqrt{\frac{1}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$ (146)
with probability at least $1-\beta/2$. We note that without the bound in $\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}$, we cannot use the trick of Equation˜112 to cancel the $\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}$ terms and get a rate in $N^{-1}$ instead of $N^{-1/2}$.
@@
Adapting the reasoning of Equations˜107, 108, 109, 110 and 111 to keep track of the new bandwidth scaling of Equation˜139, we get that Equation˜108 becomes
$\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}\penalty 10000\$$\leq\penalty 10000\ \mathbb{E}\left[h_{\lambda}(Z,Z^{\prime})^{2}\right]+\sqrt{\frac{\widetilde{\sigma}_{1}^{2}}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\penalty 10000\ \frac{1}{\lambda^{4d}}\,\mathbb{E}\left[k_{\lambda}(Z,Z^{\prime})^{2}\right]+\sqrt{\frac{1}{N\lambda^{10d}}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\penalty 10000\ \frac{1}{\lambda^{5d}}+\frac{1}{\lambda^{5d}}\sqrt{\frac{1}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\penalty 10000\ \frac{1}{\lambda^{5d}}\log\left(\frac{1}{\beta}\right)$ (147)
with probability at least $1-\beta/2$, with
[…]
Finally, following the reasoning of Equations˜107, 108, 109, 110 and 111, we obtain the quantile bound
${q}_{1-\alpha}\penalty 10000\$$\lesssim\penalty 10000\ \frac{1}{N\lambda^{5d/2}}\sqrt{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right)$ (149)
holding with probability at least $1-\beta/2$.
@@
Using Equations˜146 and 149, the type II error can be controlled as
@@
${\mathrm{Kdisc}}^{2}\penalty 10000\ \gtrsim\penalty 10000\ \frac{1}{\lambda^{5d/2}\sqrt{N}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (151)
@@
Using the expression of Equation˜140, the power guaranteeing condition of Equation˜151 becomes
$\big\|{\bm{\psi}}\big\|_{L^{2}}^{2}\penalty 10000\ \gtrsim\penalty 10000\ \big\|{\bm{\psi}}-S_{\lambda}{\bm{\psi}}\big\|_{L^{2}}^{2}-\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}+\frac{C}{\lambda^{5d/2}\sqrt{N}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (152)
The difference with the MMD case (e.g., Equation˜115) is the rate $N^{-1/2}$ instead of $N^{-1}$, the extra term $\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}$, and the rate in $\lambda$. Here, we simply bound $\left\lVert S_{\lambda}{\bm{\psi}}\right\rVert_{L^{2}}^{2}$ by 0 (instead of being able to use this term to improve the rate in $N$ as in the MMD case), we get
$\big\|{\bm{\psi}}\big\|_{L^{2}}^{2}\penalty 10000\ \gtrsim\penalty 10000\ \big\|{\bm{\psi}}-S_{\lambda}{\bm{\psi}}\big\|_{L^{2}}^{2}+\frac{C}{\lambda^{5d/2}\sqrt{N}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (153)
@@
$\int_{\mathbb{R}^{d}}\big\|\xi\big\|^{2s}_{2}\,\big\|\widehat{\bm{\psi}}(\xi)\big\|^{2}_{2}\,\mathrm{d}\xi\penalty 10000\ \leq\penalty 10000\ (2\pi)^{d}$ (154)